Tom can make 5 pizzas in 1 hour. Chris can make 7 pizzas in 2 hours. How long will it take (approximately) for Tom and Chris together to make 35 pizzas?
Answer by nerdybill(7384)

Tom can make 5 pizzas in 1 hour. Chris can make 7 pizzas in 2 hours. How long will it take (approximately) for Tom and Chris together to make 35 pizzas?
Let x = time (hours) it takes for both to make 35
then
x(5/1 + 7/2) = 35
x(10 + 7) = 70
x(17) = 70
x = 70/17
x = 4.118
or
x = 4 hours and 7 minutes